OpenStudy (anonymous):

pls pls help me... medal Use the substitution method to solve the following system of equations. 3x + y – z = 2 2x + y + 2z = 8 2x – 3y + z = 21

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Can you help me with the question pls

OpenStudy (zehanz):

If you look at the first equation, you could rewrite that as z = 3x + y -2. Use this to replace z in the other two equations and simplify. What do you get?

OpenStudy (anonymous):

but why isolating the z though @ZeHanz

OpenStudy (zehanz):

Doesn't matter which one you isolate. I just took z because it was in the first equation and I didn't have to divide to isolate it. If you isolate another, you'd do the same: replace in the two other equations.

OpenStudy (anonymous):

trust me, i tried but i couldn't get the solution. they didn't make any sense @ZeHanz

OpenStudy (anonymous):

@ZeHanz pls help me

OpenStudy (zehanz):

OK, let's see what we get with z = 3x + y -2 in eqations 2 and 3: 2x + y + 2(3x + y -2) = 8 2x – 3y + 3x + y -2 = 21 We only have x and y in these equations. Simplified: 2x+y+6x+2y-4=8 2x-3y+3x+y-2=21 8x+3y=12 5x-2y=23

OpenStudy (zehanz):

Now we can again use substitution, although this time, things are not so nice: Suppose we isolate y from the first one: 3y=12-8x, so y=4-8x/3 Suppose we isolate x from the first one: 8x=12-3y, so x=12/8-3y/8, even worse... I would uese elimination from this point (after all, we already DID use substitution): 8x+3y=12 multiply with 2 : 16x+6y=24 5x-2y=23 multiply with 3: 15x-6y=69 Add: 16x+6y=24 15x-6y=69 ---------- + 31x =93 So x=93/31=3 Now we can first find y, then z.